OpenStudy (anonymous):

Find the range of f(x)= square root 1/2x - 4+ 1 help

OpenStudy (isaiah.feynman):

The range depends on the domain.

OpenStudy (anonymous):

The range is not 1?

OpenStudy (anonymous):

I don't need a graph

OpenStudy (anonymous):

\[f(x)= \sqrt{1\div2x-4}+1\]

OpenStudy (anonymous):

I need to find the range of that equation ^

myininaya (myininaya):

\[f(x)=\sqrt{\frac{1}{2x-4}}+1 ?\]

myininaya (myininaya):

the ? after it means I'm trying to ask you is that what you meant?

myininaya (myininaya):

If so let's talk about the possible outputs... first of all looking at just the inside of the sqrt Can 1/(2x-4) ever be 0? Think about a fraction... When is a fraction 0?